# Functors on the category of finite sets revisited

### v2

@@ Abstract.
We study the structure of the category of representations of $\mathbf{FA}$, the category of finite sets and all maps, mostly working over a field of characteristic zero. This category is not semi-simple and exhibits interesting features.
We construct the simple representations explicitly, together with their projective covers. This recovers the classification of the simples given by Wiltshire-Gordon. The approach also yields a convenient set of projective generators of the category of representations of $\mathbf{FA}$ and hence a Morita equivalence. This is used to explain how to calculate the multiplicities of the composition factors of an arbitrary object, based only on its underlying $\mathbf{FB}$-representation, where $\mathbf{FB}$ is the category of finite sets and bijections.
@@ 1. Introduction
Recently, the study of representations of categories of a combinatorial nature has been of major interest. (Here a representation of an essentially small category $\mathcal{C}$ is understood to be a functor from $\mathcal{C}$ to $\mathtt{Mod}_{\mathbbm{k}}$, the category of $\mathbbm{k}$-vector spaces for a field $\mathbbm{k}$; these form a category denoted $\mathcal{F}(\mathcal{C})$.) This is especially true for categories related to $\mathbf{FA}$, the category of finite sets, including the subcategory ${\mathbf{FI}}$ of finite sets and injections, the subcategory ${\mathbf{FS}}$ of finite sets and surjections, and $\mathbf{FB}$ that of bijections (see [SS17] for example). Here our main interest is in $\mathcal{F}(\mathbf{FA})$, also referred to here as $\mathbbm{k}\mathbf{FA}$-modules.
This paper starts by establishing results that hold without supposing that $\mathbbm{k}$ has characteristic zero. One has the family of standard projective generators $P^{\mathbf{FA}}_{\mathbf{n}}$, for $n\in\mathbb{N}$, where $P^{\mathbf{FA}}_{\mathbf{n}}(X)=\mathbbm{k}[X]^{\otimes n}$, for a finite set $X$; $\mathbf{n}$ denotes the set $\{1,\ldots,n\}$. The functor $\overline{P}$ is defined by the short exact sequence
@@
The functors $\overline{P}^{\otimes n}$ are not projective but are not far from being so. Indeed, the functor $\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$ is projective (see Corollary 3.10) and this sits in a short exact sequence
$0\rightarrow\overline{P}^{\otimes n+1}\rightarrow\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{P}^{\otimes n}\rightarrow 0$
The family $\{\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}\mid n\in\mathbb{N}\}\amalg\{P^{\mathbf{FA}}_{\mathbf{0}}=\mathbbm{k}\}$ gives a much more efficient set of projective generators of $\mathcal{F}(\mathbf{FA})$, as stated in Corollary 3.11.
@@
where ${\mathbf{FS}}$ is the wide subcategory of $\mathbf{FA}$ given by the surjective maps; this surjection is defined by sending non-surjective maps to zero. Using an analysis of the projectives $\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$, one has (see Theorem 4.8):
Theorem 1.
@@
induced by $\overline{P}^{\otimes n}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{n}}$ identifies with (1.2).
For this, it is not difficult to obtain the induced injection $\mathbbm{k}{\mathbf{FS}}(-,\mathbf{n})\hookrightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})$; less immediate is that this is an isomorphism, since $\overline{P}^{\otimes n}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{n}}$ does not admit a retract.
We then turn to the case where $\mathbbm{k}$ has characteristic zero. In this case, the simple $\mathbbm{k}\mathbf{FA}$-modules were classified by Wiltshire-Gordon [Wil14]. In [SST22], Sam, Snowden and Tosteson showed that the category of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules is equivalent to the category of polynomial representations of the Witt Lie algebra. Using this, they observed that Wiltshire-Gordon’s simple modules correspond to a family of representations of the Witt Lie algebra constructed by Feigin and Shoikhet [FS07]. Part of the paper revisits this classification using our more efficient set of projective generators.
@@
that sends non-injective maps to zero, analogous to (1.2). This is a morphism of $\mathbbm{k}\mathfrak{S}_{n}^{\mathrm{op}}\otimes\mathbbm{k}\mathbf{FA}$-modules. Now, for $n$ a positive integer and $\lambda\vdash n$ a partition such that $\lambda\neq(1^{n})$, with associated simple $\mathfrak{S}_{n}$-module $S_{\lambda}$ (which is thus not isomorphic to the sign representation $\mathsf{sgn}_{n}$), one has the $\mathbbm{k}\mathbf{FA}$-module:
@@
Here one uses the convention that $\tilde{C}_{(1^{0})}=\mathbbm{k}_{\mathbf{0}}$, the functor supported on $\emptyset$ with value $\mathbbm{k}$, and $\tilde{C}_{(1)}=\overline{\mathbbm{k}}$. There is a surjection $\Lambda^{n+1}(P^{\mathbf{FA}}_{\mathbf{1}})\twoheadrightarrow\Lambda^{n}(\overline{P})$.
@@
This recovers Wiltshire-Gordon’s classification [Wil14], giving an explicit description of the simples and also of their projective covers. This construction of the $\tilde{C}_{\lambda}$ also establishes the relation with the work of Feigin and Shoikhet [FS07], which can be interpreted (via the results of [SST22]) as being equivalent to the analysis of $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ as a $\mathbbm{k}\mathbf{FA}$-module. The latter is almost semi-simple:
Corollary 3.
@@
Now, whilst $\overline{P}^{\otimes n}$ is not projective, over a field of characteristic zero, it is very close to being so. More precisely, writing $\mathbb{P}_{n}$ for the projective cover of $\overline{P}^{\otimes n}$, one has a $\mathfrak{S}_{n}^{\mathrm{op}}$-equivariant exact sequence
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\otimes\mathcal{S}(k-1),$
@@
    Theorem 10.6 determines $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})$, thus giving an answer to the ‘urgent question’ posed by Wiltshire-Gordon in [Wil14, Section 5.2] asking how to calculate the composition factors of the projectives $S_{\lambda}\otimes_{\mathfrak{S}_{n}}P^{\mathbf{FA}}_{\mathbf{n}}$, for partitions $\lambda\vdash n$.
@@
    Corollary 10.9 determines $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\mathbb{P}_{n})$, an essential step towards applying the Morita equivalence result, Theorem 4.
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(\mathbf{n},\mathbf{k}))\otimes[\mathsf{sgn}_{k-1}].$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\sum_{t\in\mathbb{N}}(-1)^{t}[\mathsf{sgn}_{k+t}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
Theorem 7 recovers, by very different methods, one of the results of [PV23] (see also [Pow24a]). The relationship between the methods used here and in those works is outlined in Section 4.4.
Organization: The paper is presented in three Parts, separating out the main themes. Part I provides background and establishes the results that do not require working over a field of characteristic zero. Part II gives the classification of the simple $\mathbbm{k}\mathbf{FA}$-modules and their projective covers, working over a field of characteristic zero; this yields the Morita equivalence. Part III shows how to carry out associated basic calculations, including how to calculate the composition factors of a $\mathbbm{k}\mathbf{FA}$-module.
Acknowledgements: The precursor of this research was the author’s work with Christine Vespa [PV23], which took a very different approach (and worked over a general commutative ring $\mathbbm{k}$). The author is very grateful to Christine Vespa for her ongoing interest.
The author also thanks Vladimir Dotsenko for an invitation to Strasbourg in January 2024 (financed by his IUF) and for useful discussions - during which the work of Feigin and Shoikhet appeared from a different viewpoint! Some of the research underlying this paper was carried out during that visit.
Part I Arbitrary characteristic
@@ 2. Background
The purpose of this section is to recall some generalities on the structure of the category of $\mathbbm{k}\mathcal{C}$-modules, where $\mathcal{C}$ is an essentially small category. Then we specialize to the cases of interest, related to the category $\mathbf{FA}$ of finite sets. This serves to introduce notation and some basic constructions. Throughout, $\mathbbm{k}$ is a field.
@@ 2.1. Basics
For $\mathcal{C}$ an essentially small category, $\mathcal{F}(\mathcal{C})$ denotes the category of functors from $\mathcal{C}$ to $\mathtt{Mod}_{\mathbbm{k}}$, the category of $\mathbbm{k}$-vector spaces. It is equivalent to the category of $\mathbbm{k}$-linear functors from $\mathbbm{k}\mathcal{C}$ (the $\mathbbm{k}$-linearization of $\mathcal{C}$) to $\mathbbm{k}$-vector spaces. Such functors will be referred to as $\mathbbm{k}\mathcal{C}$-modules.
The category $\mathcal{F}(\mathcal{C})$ is abelian and is equipped with the tensor product $\otimes$ defined value-wise: for functors $F$, $G$ in $\mathcal{F}(\mathcal{C})$, $(F\otimes G)(X):=F(X)\otimes G(X)$ (forming $\otimes:=\otimes_{\mathbbm{k}}$ in $\mathtt{Mod}_{\mathbbm{k}}$).
Vector space duality induces an exact functor $D:\mathcal{F}(\mathcal{C})^{\mathrm{op}}\rightarrow\mathcal{F}(\mathcal{C}^{\mathrm{op}})$ where, for a $\mathbbm{k}\mathcal{C}$-module $F$, the dual $DF$ is defined by $DF(X):=\mathrm{Hom}_{\mathtt{Mod}_{\mathbbm{k}}}(F(X),\mathbbm{k})$.
@@
The tensor product $\otimes_{\mathfrak{S}_{n}}$ for representations of the symmetric group $\mathfrak{S}_{n}$ has the global form $\otimes_{\mathbf{FB}}:\mathcal{F}(\mathbf{FB}^{\mathrm{op}})\times\mathcal{F}(\mathbf{FB})\rightarrow\mathtt{Mod}_{\mathbbm{k}}$. It is given for a $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-module $X$ and a $\mathbbm{k}\mathbf{FB}$-module $M$ by:
@@
$\mathbbm{k}\mathbf{FA}\cong\mathbbm{k}{\mathbf{FI}}\otimes_{\mathbf{FB}}\mathbbm{k}{\mathbf{FS}}$
(the variance dictates how $\otimes_{\mathbf{FB}}$ is formed).
@@
The functor $\mathbbm{k}{\mathbf{FI}}\otimes_{\mathbf{FB}}-:\mathcal{F}(\mathbf{FB})\rightarrow\mathcal{F}(\mathbf{FB})$ is naturally equivalent to the functor $\mathsf{triv}\odot-:\mathcal{F}(\mathbf{FB})\rightarrow\mathcal{F}(\mathbf{FB})$.
@@
Now, for $s\in\mathbb{N}$ (that we may assume to be at most $n$), $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{n})$ is isomorphic to the $\mathbbm{k}(\mathfrak{S}_{n}\times\mathfrak{S}_{s}^{\mathrm{op}})$-module $\mathbbm{k}\mathfrak{S}_{n}/\mathfrak{S}_{n-s}$, with $\mathfrak{S}_{s}$ acting on the right via the residual right regular action (use the inclusion of Young subgroup $\mathfrak{S}_{n-s}\times\mathfrak{S}_{s}\subset\mathfrak{S}_{n}$) and $\mathfrak{S}_{n}$ acts via the left regular action. Hence one sees that $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{n})\otimes_{\mathfrak{S}_{s}}M(s)$ is naturally isomorphic to $\mathsf{triv}_{n-s}\odot M(s)$. Summing over $s$ gives the result. ∎

### v3

@@ Abstract.
We study the structure of the category of representations of $\mathbf{FA}$, the category of finite sets and all maps, working over a unital commutative ring. We first exhibit a convenient set of projective generators by exploiting Möbius inversion and establish some general structural results.
We then specialize to working over a field of characteristic zero. We construct the simple representations explicitly, together with their projective covers. This recovers the classification of the simples given by Wiltshire-Gordon.
Hence we obtain a more efficient set of projective generators of the category of representations of $\mathbf{FA}$ and the associated Morita equivalence. This is used to explain how to calculate the multiplicities of the composition factors of an arbitrary object, based only on its underlying $\mathbf{FB}$-representation, where $\mathbf{FB}$ is the category of finite sets and bijections.
@@
Functors on finite sets; simple functors; projective covers; Morita equivalence
2020 Mathematics Subject Classification:
18A25
@@ 1. Introduction
Recently, the study of representations of categories of a combinatorial nature has been of major interest. (Here a representation of an essentially small category $\mathcal{C}$ is understood to be a functor from $\mathcal{C}$ to $\mathtt{Mod}_{\mathbbm{k}}$, the category of $\mathbbm{k}$-modules over a unital commutative ring $\mathbbm{k}$; these form a category denoted $\mathcal{F}(\mathcal{C})$.) This is especially true for categories related to $\mathbf{FA}$, the category of finite sets, including the subcategory $\mathbf{FI}$} of finite sets and injections, the subcategory $\mathbf{FS}$} of finite sets and surjections, and $\mathbf{FB}$ that of bijections (see [8] for example). Here our main interest is in $\mathcal{F}(\mathbf{FA})$, also referred to here as $\mathbbm{k}\mathbf{FA}$-modules.
This paper starts by establishing some results that hold without supposing that $\mathbbm{k}$ is a field. One has the family of standard projective generators $P^{\mathbf{FA}}_{\mathbf{n}}$, for $n\in\mathbb{N}$, where $P^{\mathbf{FA}}_{\mathbf{n}}(X)=\mathbbm{k}[X]^{\otimes n}$, for a finite set $X$; $\mathbf{n}$ denotes the set $\{1,\ldots,n\}$. The functor $\overline{P}$ is defined by the short exact sequence
@@
The functors $\overline{P}^{\otimes n}$ are not projective but are not far from being so. Indeed, the functor $\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$ is projective (see Proposition 3.10) and this sits in a short exact sequence
$0\rightarrow\overline{P}^{\otimes n+1}\rightarrow\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow\overline{P}^{\otimes n}\rightarrow 0.$
Moreover, the family $\{\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}\mid n\in\mathbb{N}\}\amalg\{P^{\mathbf{FA}}_{\mathbf{0}}=\mathbbm{k}\}$ gives a more efficient set of projective generators of $\mathcal{F}(\mathbf{FA})$. This is proved in Section 4 by using the general framework of Möbius inversion, thus giving a conceptual explanation of this result (see Corollary 4.8).
In Section 4, for $(X,x)$ a finite pointed set, we work with an explicit direct summand $\mathbf{Q}_{\emptyset}^{X,x}$ of $P^{\mathbf{FA}}_{X}$, which is shown to be isomorphic to $P^{\mathbf{FA}}_{\mathbf{1}}\otimes\overline{P}^{\otimes X\backslash\{x\}}$. The explicit construction allows some control of morphisms between such objects; in particular we prove the fundamental result:
Theorem 1.1.
(Theorem 4.19.) For finite pointed sets $(X,x)$ and $(Y,y)$, there are natural isomorphisms:
[…]
This then yields some information on morphisms between functors of the form $\overline{P}^{\otimes n}$:
Corollary 1.2.
(Corollary 4.21.) For $s,t\in\mathbb{N}$, if $s>t$ then
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes s},\overline{P}^{\otimes t})=0.$
For $s=t$, the place permutation action on $\overline{P}^{\otimes t}$ induces an isomorphism of $\mathbbm{k}$-algebras:
$\mathbbm{k}\mathfrak{S}_{t}\cong\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes t}).$
One consequence of these results is that, when $\mathbbm{k}$ is a field, the indecomposable projectives in $\mathcal{F}(\mathbf{FA})$ admit (finite length) composition series (see Corollary 4.33).
@@
where $\mathbf{FS}$} is the wide subcategory of $\mathbf{FA}$ given by the surjective maps; this surjection is defined by sending non-surjective maps to zero. Using the analysis of the projectives $\overline{P}^{\otimes n}\otimes P^{\mathbf{FA}}_{\mathbf{1}}$, one has (see Corollary 4.25):
Theorem 1.3.
@@
induced by $\overline{P}^{\otimes n}\hookrightarrow P^{\mathbf{FA}}_{\mathbf{n}}$ identifies with (1.2). In particular, there is an isomorphism $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\cong\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n})$.
We then turn to the case where $\mathbbm{k}$ is a field of characteristic zero. In this case, the simple $\mathbbm{k}\mathbf{FA}$-modules were classified by Wiltshire-Gordon [10]. In [9], Sam, Snowden and Tosteson showed that the category of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules is equivalent to the category of polynomial representations of the Witt Lie algebra. Using this, they observed that Wiltshire-Gordon’s simple modules correspond to a family of representations of the Witt Lie algebra constructed by Feigin and Shoikhet [2]. Part of the paper revisits this classification using our more efficient set of projective generators.
@@
sending non-injective maps to zero, analogous to (1.2). This is a morphism of $\mathbbm{k}\mathfrak{S}_{n}^{\mathrm{op}}\otimes\mathbbm{k}\mathbf{FA}$-modules. Now, for $n$ a positive integer and $\lambda\vdash n$ a partition such that $\lambda\neq(1^{n})$, with associated simple $\mathfrak{S}_{n}$-module $S_{\lambda}$ (which is thus not isomorphic to the sign representation $\mathsf{sgn}_{n}$), one has the $\mathbbm{k}\mathbf{FA}$-module:
@@
Here we use the convention that $\tilde{C}_{(1^{0})}=\mathbbm{k}_{\mathbf{0}}$, the functor supported on $\emptyset$ with value $\mathbbm{k}$, and $\tilde{C}_{(1)}=\overline{\mathbbm{k}}$. There is a surjection $\Lambda^{n+1}(P^{\mathbf{FA}}_{\mathbf{1}})\twoheadrightarrow\Lambda^{n}(\overline{P})$.
@@
    if $\lambda\neq(1^{n})$, $S_{\lambda}(\overline{P})\rightarrow\tilde{C}_{\lambda}$ is the projective cover of $\tilde{C}_{\lambda}$;
@@
This recovers Wiltshire-Gordon’s classification [10], giving an explicit description of the simples and also of their projective covers. This construction of the $\tilde{C}_{\lambda}$ also establishes the relation with the work of Feigin and Shoikhet [2], which can be interpreted (via the results of [9]) as being equivalent to the analysis of $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ as a $\mathbbm{k}\mathbf{FA}$-module. The latter is almost semi-simple:
Corollary 1.5.
@@
Now, whilst $\overline{P}^{\otimes n}$ is not projective, it is very close to being so. More precisely, writing $\mathbb{P}_{n}$ for the projective cover of $\overline{P}^{\otimes n}$, one has a $\mathfrak{S}_{n}^{\mathrm{op}}$-equivariant exact sequence
@@
Remark 1.6.
For $n\in\mathbb{N}$, there is a counterpart for the structure of the $\mathbbm{k}\mathbf{FA}$-module $D\mathbbm{k}{\mathbf{FS}}(-,\mathbf{n})$, where $D$ is the duality functor $\mathcal{F}(\mathbf{FA}^{\mathrm{op}})^{\mathrm{op}}\rightarrow\mathcal{F}(\mathbf{FA})$ induced by vector space duality. Namely, Theorem 1.3 together with the fact that $\overline{P}^{\otimes n}$ is ‘almost’ a direct summand of $P^{\mathbf{FA}}_{\mathbf{n}}$ imply that $D\mathbbm{k}{\mathbf{FS}}(-,\mathbf{n})$ is ‘almost’ an injective $\mathbbm{k}\mathbf{FA}$-module. The failure to be injective is encoded by the dual of the short exact sequence given in Proposition 10.3. (These observations are not developed further in this paper.)
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},F)]=[\overline{F}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}F(\mathbf{k}))\otimes\mathcal{S}(k-1),$
@@
    Theorem 10.6 determines $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})$, thus giving an answer to the ‘urgent question’ posed by Wiltshire-Gordon in [10, Section 5.2] asking how to calculate the composition factors of the projectives $S_{\lambda}\otimes_{\mathfrak{S}_{n}}P^{\mathbf{FA}}_{\mathbf{n}}$, for partitions $\lambda\vdash n$.
@@
    Corollary 10.9 determines $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\mathbb{P}_{n})$, an essential step towards applying the Morita equivalence result, Theorem 1.7.
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(\mathbf{n},\mathbf{k}))\otimes[\mathsf{sgn}_{k-1}].$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\ +\ \sum_{k\geq 1}\mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\ +\ \sum_{k\geq 1}\sum_{t\in\mathbb{N}}(-1)^{t}[\mathsf{sgn}_{k+t}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
Theorem 1.10 recovers, by very different methods, one of the results of [7] (see also [5]).
Organization: The paper is presented in three Parts, separating out the main themes. Part I provides background and establishes some general results that hold over any unital commutative ring. Part II gives the classification of the simple $\mathbbm{k}\mathbf{FA}$-modules and their projective covers, working over a field of characteristic zero; this yields the Morita equivalence. Part III shows how to carry out associated basic calculations, including how to calculate the composition factors of a $\mathbbm{k}\mathbf{FA}$-module.
Acknowledgements: The precursor of this research was the author’s work with Christine Vespa [7], which took a very different approach. The author is very grateful for her ongoing interest.
The author also thanks Vladimir Dotsenko for an invitation to Strasbourg in January 2024 (financed by his IUF) and for useful discussions - during which the work of Feigin and Shoikhet appeared from a different viewpoint. Some of the research underlying this paper was carried out during that visit.
The author is also immensely grateful to the referee, both for their careful reading of the manuscript and for their suggestion that Möbius inversion gives a conceptual approach to the general results of Part I, also strengthening some results, for which they provided details. This lead to the current, less ad hoc presentation of Section 4 and also lead the author to only impose the hypothesis that $\mathbbm{k}$ is a field when necessary.
Part I Working over a unital commutative ring
@@ 2. Background
The purpose of this section is to recall some generalities on the structure of the category of $\mathbbm{k}\mathcal{C}$-modules, where $\mathcal{C}$ is an essentially small category. Then we specialize to the cases of interest, related to the category $\mathbf{FA}$ of finite sets. This serves to introduce notation and some basic constructions. Unless otherwise stated, $\mathbbm{k}$ is a unital commutative ring.
@@ 2.1. Basics
For $\mathcal{C}$ an essentially small category, $\mathcal{F}(\mathcal{C})$ denotes the category of functors from $\mathcal{C}$ to $\mathtt{Mod}_{\mathbbm{k}}$, the category of $\mathbbm{k}$-modules. It is equivalent to the category of $\mathbbm{k}$-linear functors from $\mathbbm{k}\mathcal{C}$ (the $\mathbbm{k}$-linearization of $\mathcal{C}$) to $\mathtt{Mod}_{\mathbbm{k}}$. Such functors will be referred to as $\mathbbm{k}\mathcal{C}$-modules.
The category $\mathcal{F}(\mathcal{C})$ is abelian: a complex in $\mathcal{F}(\mathcal{C})$ is exact if and only if, for each $X\in\mathrm{Ob}\hskip 2.0pt\mathcal{C}$, the complex in $\mathtt{Mod}_{\mathbbm{k}}$ obtained by evaluating on $X$ is exact. This category is equipped with the tensor product $\otimes$ defined value-wise: for functors $F$, $G$ in $\mathcal{F}(\mathcal{C})$, $(F\otimes G)(X):=F(X)\otimes G(X)$ (forming $\otimes:=\otimes_{\mathbbm{k}}$ in $\mathtt{Mod}_{\mathbbm{k}}$).
@@
The tensor product $\otimes_{\mathfrak{S}_{n}}$ for representations of the symmetric group $\mathfrak{S}_{n}$ has a ‘global’ form $\otimes_{\mathbf{FB}}:\mathcal{F}(\mathbf{FB}^{\mathrm{op}})\times\mathcal{F}(\mathbf{FB})\rightarrow\mathtt{Mod}_{\mathbbm{k}}$. It is given for a $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-module $X$ and a $\mathbbm{k}\mathbf{FB}$-module $M$ by:
@@
$\mathbbm{k}\mathbf{FA}\cong\mathbbm{k}{\mathbf{FI}}\otimes_{\mathbf{FB}}\mathbbm{k}{\mathbf{FS}},$
where the variance dictates how $\otimes_{\mathbf{FB}}$ is formed.
@@
The composite of the functor $\mathbbm{k}{\mathbf{FI}}\otimes_{\mathbf{FB}}-:\mathcal{F}(\mathbf{FB})\rightarrow\mathcal{F}({\mathbf{FI}})$ with the restriction $\mathcal{F}({\mathbf{FI}})\rightarrow\mathcal{F}(\mathbf{FB})$ induced by the inclusion $\mathbf{FB}\subset{\mathbf{FI}}$ is naturally equivalent to the functor $\mathsf{triv}\odot-:\mathcal{F}(\mathbf{FB})\rightarrow\mathcal{F}(\mathbf{FB})$.
@@
Now, for $s\in\mathbb{N}$ (that we may assume to be at most $n$), $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{n})$ is isomorphic to the $\mathbbm{k}(\mathfrak{S}_{n}\times\mathfrak{S}_{s}^{\mathrm{op}})$-module $\mathbbm{k}\mathfrak{S}_{n}/\mathfrak{S}_{n-s}$, with $\mathfrak{S}_{s}$ acting on the right via the residual right regular action (use the inclusion of Young subgroup $\mathfrak{S}_{n-s}\times\mathfrak{S}_{s}\subset\mathfrak{S}_{n}$) and the left $\mathfrak{S}_{n}$-action is the usual one. Hence one sees that $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{n})\otimes_{\mathfrak{S}_{s}}M(s)$ is naturally isomorphic to $\mathsf{triv}_{n-s}\odot M(s)$. Summing over $s$ gives the result. ∎